# Optimizing Byte-level Representation for End-to-end ASR

### v1

@@ Abstract
keywords:
byte-level representation, speech recognition
@@ 1 Introduction
End-to-end (E2E) neural networks are flexible and accurate models for multilingual automatic speech recognition (ASR). The output of such a multilingual model is often unions of characters or subwords of the supported languages. However, as the number of languages increases, the size of the output layer increases, which can negatively affect compute, memory usage and asset size. This problem is more prominent when the system supports languages that have large character sets, such as Chinese, Japanese and Korean (CJK). To tackle this problem, previous work proposed the use of byte level representation [1, 2]. By using UTF-8 [3] codewords as the underlying base tokens, the output vocabulary is no longer constrained by the character sets of each language, allowing developers to choose a vocabulary size based on compute, and memory constraints. One well-known multilingual ASR system that uses UTF-8 subwords is Whisper [4].
@@
For the first point, we think the representation should be optimized for the target machine learning task. The design of the representation should boost accuracy, and be data driven. For the second point, we think the design should take advantage of all available information. For example, if the representation is designed for ASR, it should incorporate the information from both text and audio. Ideally, the framework should be flexible enough to incorporate any available information. For ASR, it means it should be able to take lexicon, phonemes or other useful information into account. For the last property, we believe the representation should provide an error correction mechanism to handle the cases when the model produces an invalid sequence, and the recovery should aim for minimal error.
@@ 3.2 Formulation as auto-encoding optimization problem
Figure 1: Overview of the proposed auto-encoder
@@ 3.5 Comparing UTF-8 and VQ based representation
While UTF-8 and our proposed VQ based representations are both byte-level representations, they have a few key differences. First, UTF-8 is a variable length prefix code, but our proposed VQ representation is a fixed length code. Second, VQ representation can be optimized for a specific machine learning task while UTF-8 is fixed and not optimized for machine learning. Third, UTF-8 recovers invalid sequences with the sequences with most characters, while VQ’s label decoder would try to find the most likely output given any input sequence. Finally, as a learned representation, it is possible to have collision, i.e. different characters can be mapped to the same byte sequences, which would not be the case for UTF-8.
@@ 4 Experimental results
We evaluate our approach through our proprietary English and Mandarin dictation tasks. We build bilingual English and Mandarin ASR models using a mixture of monolingual English and Mandarin supervised data, which is randomly sampled and anonymized. The English training set consists of 10k hours of data while the Mandarin training set has 14k hours of data. Our training sets cover domains like dictation and assistant use cases. For test sets, we use the dictation test sets for English and Mandarin, which are randomly sampled and anonymized. The English test set consists of 27 hours of data and the Mandarin test set has 13 hours of data. Both train and test sets cover platforms like smart phones, tablets and laptops. When we report accuracy numbers on these test sets, we report word error rate (WER) for English and character error rate (CER) for Mandarin. We use the term token error rate (TER) when we discuss the error rates of both languages.
@@
Table 1: TER(%) of CTC-AED model using character based, UTF-8 and our proposed VQ based representation. The VQ based representation has three codebooks and each codebook has 256 embeddings ($N=3$, $M=256$). Both UTF-8 and VQ use BPE to create subword output, and each column uses different number of subwords in the ASR output.
@@ 4.1 Ablation study
Table 3 shows the results of applying different weights to the cross entropy loss with respect to the acoustic encoder (the second term of Equation 3). When the weight is zero, it means the acoustic encoder would not contribute to representation learning. In this case, the auto-encoder would ignore acoustic information and the acoustic encoder would fit to the latent variables decided by the label encoder. The results in Table 3 show that acoustic information helps.
@@ 5 Conclusions and future work
In this paper, we propose an algorithm to optimize a byte-level representation for ASR and compare it with UTF-8 representation. Our proposed representation can be optimized with both audio and text data, and the error correction mechanism is optimized for accuracy. On our English/Mandarin dictation test sets, our auto-encoder and VQ based approach can achieve 5% relative reduction in TER.
While our proposed approach has accuracy advantages, UTF-8 has some desirable properties, such as, being a prefix code, not having a collision issue. In addition, the variable length design is more flexible, though ideally, the length would be correlated with token frequency. In the future, we will explore VQ methods that allow variable length. If code length can be optimized jointly, we would be able to bypass BPE and use that to generate target subwords directly. While as a machine learning approach, it cannot guarantee no collision in the learned representation, we will look into whether we can derive a prefix code from the auto-encoder. This would also make encoding and decoding more efficient. Last, but not least, we will also look into using other side information, such as a phonemic lexicon, to improve the representation.

### v2

@@ Abstract
Index Terms—  byte-level representation, speech recognition
@@ 1 Introduction
End-to-end (E2E) neural networks are flexible and accurate models for multilingual automatic speech recognition (ASR). The output of such a multilingual model is often unions of characters or subwords of the supported languages. However, as the number of languages increases, the size of the output layer increases, which can negatively affect compute, memory usage and asset size. This problem is more prominent when the system supports languages that have large character sets, such as Chinese, Japanese and Korean (CJK). To tackle this problem, previous work proposed the use of byte level representation for E2E ASR [1, 2]. By using UTF-8 [3] codewords as the underlying base tokens, the output vocabulary is no longer constrained by the character sets of each language, allowing developers to choose a vocabulary size based on compute, and memory constraints. One well-known multilingual ASR system that uses UTF-8 subwords is Whisper [4].
@@
For the first point, we think the representation should be optimized for the target machine learning task. The design should boost accuracy, and be data driven. For the second point, we think the design should consider all available information. For example, if the representation is designed for ASR, it should incorporate the information from both text and audio. Ideally, the framework should be flexible enough to incorporate any available information. For ASR, it means it should be able to take lexicon, phonemes or other useful information into account. For the last property, we believe the representation should provide an error correction mechanism to handle the cases when the model produces an invalid sequence, and the recovery should aim for minimal error.
@@ 3.2 Formulation as auto-encoding optimization problem
Fig. 1: Overview of the proposed auto-encoder. The label tokens could be words/subwords/characters. In our experiments, the label tokens are characters.
@@ 3.5 Comparing UTF-8 and VQ based representation
While UTF-8 and our proposed VQ based representations are both byte-level representations, they have a few key differences. First, UTF-8 is a variable length prefix code, but our proposed VQ representation is a fixed length code. Second, VQ representation can be optimized for a specific machine learning task while UTF-8 is fixed and not optimized for machine learning. Third, UTF-8 recovers invalid sequences with the sequences with most characters [1], while VQ’s label decoder would try to find the most likely output given any input sequence. Finally, as a learned representation, it is possible to have collision, i.e. different characters can be mapped to the same byte sequences, which is not the case for UTF-8.
@@ 4 Experimental results
We evaluate our approach through our proprietary English and Mandarin dictation tasks. Since our proposed approach is new, we would like to focus on simpler cases where the ASR system only handles two languages. We build these bilingual English and Mandarin ASR models using a mixture of monolingual English and Mandarin supervised data, which is randomly sampled and anonymized. The English training set consists of 10k hours of data while the Mandarin training set has 14k hours of data. Our training sets cover domains like dictation and assistant use cases. For test sets, we use the dictation test sets for English and Mandarin, which are randomly sampled and anonymized. The English test set consists of 27 hours of data and the Mandarin test set has 13 hours of data. Both train and test sets cover platforms like smart phones, tablets and laptops. When we report accuracy numbers on these test sets, we report word error rate (WER) for English and character error rate (CER) for Mandarin. We use the term token error rate (TER) when we discuss the error rates of both languages.
@@
Table 1: TER(%) of CTC-AED model using character based, UTF-8 and our proposed VQ based representation. The VQ based representation has three codebooks and each codebook has 256 embeddings ($N=3$, $M=256$). Both UTF-8 and VQ use BPE to create subword output with their corresponding error correction algorithms. Each column uses different number of subwords in the ASR output.
@@ 4.1 Ablation study
Table 3 shows the results of applying different weights to the cross entropy loss with respect to the acoustic encoder (the second term of Equation 3). When the weight is zero, it means the acoustic encoder would not contribute to representation learning. In this case, the auto-encoder would ignore acoustic information and the acoustic encoder would fit to the latent variables decided by the label encoder. The results in Table 3 show that acoustic information helps, and this supports the core concept of this paper, where optimizing a representation with all available information is beneficial.
@@ 5 Conclusions and future work
In this paper, we propose a novel algorithm to optimize a byte-level representation for ASR and compare it with UTF-8 representation. Our proposed representation can be optimized with both audio and text data, and the error correction mechanism is optimized for accuracy. On our English/Mandarin dictation test sets, our auto-encoder and VQ based approach can achieve 5% relative reduction in TER.
As this proposed algorithm is new, we focus on simpler bilingual ASR in this paper. To learn a representation that can cover all languages, we believe there are multiple issues that need to be addressed, like the index collapse issue [16]. The index collapse issue is well studied within the machine learning community [17] and we will explore some of the ideas in that space. We hope that as the algorithm becomes more mature, we can learn a universal byte-level representation for all languages that functions like UTF-8.
Comparing UTF-8 and our proposed algorithm, while our approach has accuracy advantages, UTF-8 has some desirable properties, such as, being a prefix code, not having a collision issue. In addition, the variable length design is more flexible, though ideally, the length would be correlated with token frequency. In the future, we will explore VQ methods that allow variable length. If code length can be optimized jointly, we would be able to bypass BPE and use that to generate target subwords directly. While as a machine learning approach, it cannot guarantee no collision in the learned representation, we will look into whether we can derive a prefix code from the auto-encoder. This would also make encoding and decoding more efficient. Then, we will also look into using other side information, such as a phonemic lexicon, to improve the representation.